The Older Futhark, 10pt

Letter to type in ’Real Letter’ Rune
F  F  F
U  U  U
X  TH  X
A  A  A
R  R  R
K  K  K
G  G  G
W  W  W
H  H  H
N  N  N
I  I  I
J  J  J
Y  EI  Y
P  P  P
Z  Z  T
S  S  S
T  T  T
B  B  B
E  E  E
M  M  M
L  L  L
Q  ING  Q
D  D  D
O  O  O
:  :  :

FUTHARK, 10pt FUXARK FUTHARK, 20pt

FUXARK

FUTHARK, 30pt

FUXARK

FUTHARK, 40pt

FUXARK

FUTHARK, 50pt

FUXARK